BIOMEDICAL.: Cholesterol. An experimental drug lowers patient's blood serum cholesterol at the rate of t square root of 25 - t^2 units per day, where t is the number of days since the drug was administered. (0 is less than or equal to t less than or equal to 5). Find the total change during the first 3 days.
1
Expert's answer
2014-04-25T10:11:27-0400
We need to find the change during the first three days. It equals the sum of changes on day№0 + day№1+day№2+day№3:
0*sqrt(25 - 0^2)+1*sqrt(25 - 1^2)+2*sqrt(25 - 2^2)+3*sqrt(25-3^2) = 0 + sqrt(24) + 2*sqrt(21) + 3 * sqrt(16) = 2*sqrt(6) + 2* sqrt(21) + 12 = (approximately) 26.064
Answer: 26.064